Two similar bar magnets P and Q, each of magnetic moment M, are taken. If P is cut along its axial line and Q is cut along its equatorial line, all the four pieces have magnetic moment=? And also what about their pole strengths?
 

Grade:12th Pass

2 Answers

Arun
4 years ago
 

 If pole strength, magnetic moment and length of each part are m',M' and L' respectivley then ,

 

m'=m/2

m'=m

L'=L

L'=L/2

M'=M/2

Khimraj
4 years ago

m'=m/2

m'=m

L'=L

L'=L/2

M'=M/2

…...............................................................................